Use factoring to solve the following quadratic equation
x(5x+2) = 2x^ 2 -6x+16

Answer by LinnW(1048)

x(5x+2) = 2x^ 2 -6x+16
5x^2 + 2x = 2x^2 -6x + 16
add -2x^2 to each side
3x^2 + 2x = -6x + 16
add 6x to each side
3x^2 + 8x = 16
add -16 to each side
3x^2 + 8x - 16 = 0
product of outer coefficients is -48
(12)(-4)= -48
3x^2 +12x -4x - 16 = 0
3x(x + 4) -4(x + 4) = 0
(3x - 4)(x + 4) = 0
3x - 4 = 0
3x = 4
x = 4/3
x + 4 = 0
x = -4
So solutions are 4/3 and -4